OpenStudy (anonymous):

Linda received a credit card through the mail. The interest rate charged by the credit card company is 24.9%. If she maintains an average balance of $2,400 owed on the card, how much will she have paid in interest after one year?

OpenStudy (anonymous):

interest rate = 24.9 % p.a hence = 24.9 /12 = 2.075% p.m Assuming Linda pay's off her owed amount at every month end, she will have to pay interest @ 2.075% p.m on $ 2400 = 2400 * 0.02075 = 49.8$ p.m = 49.8 * 12 = 597.6 $ p.a

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

No Problem :)